# Enhancing Function-Calling Capabilities in LLMs: Strategies for Prompt Formats, Data Integration, and Multilingual Translation

### v1

@@ 1 Introduction
Figure 1: An illustration of prompt templates used for function calling and instruction following. During training, LLMs are provided with conditional prompts (shown on the left) and tasked with generating appropriate text completions (shown on the right). The special tokens <s>, <|im_start|>, <|im_end|>, <|answer|>, and <|use_tool|> are each represented as a single token after tokenization. More details are described in Section 3.1.
@@ 3.1 Prompt Templates for Function Calling and Instruction Following
We employ a tuning-based approach to enable both function-calling and instruction-following capabilities in our LLMs. This involves fine-tuning pretrained base models using prompt templates based on the Chat Markup Language (ChatML), a widely adopted format introduced by OpenAI. Two main strategies for incorporating function descriptions into prompts are explored: (1) introducing a dedicated role, such as tools, to represent function descriptions in JSON format (Figure 1(b)); and (2) embedding function descriptions alongside usage instructions within the system role (Figure 1(c)). In the latter strategy, both instruction-following and function-calling are guided by the system prompt.
During training, the LLMs are provided with conditional prompts as described above and are tasked with generating appropriate text completions. Based on the context, the fine-tuned model dynamically decides whether to respond directly or invoke functions. If no relevant functions are available, the model directly answers the query (Figure 1(d)). Otherwise, it generates structured function calls, such as [func1(arg1=val1, arg2=val2, ...), func2(...), ...], to address the query (Figure 1(f)).
To support these capabilities, we curated a diverse dataset comprising instruction-following and function-calling examples for fine-tuning. Details of the fine-tuning methodology are provided in Section 4.1. Furthermore, we evaluate the impact of different strategies for presenting function descriptions and demonstrate that integrating instruction-following data significantly enhances the model’s function-calling performance, as discussed in Section 4.2.
@@ 3.2 Decision Token
To address this, we propose the novel Decision Token. LLMs generate responses through next token prediction, where each step involves a classification task to select the next token. The Decision Token concept involves fine-tuning LLMs to perform a classification task before providing any detailed response—specifically, deciding whether to answer directly or use function calling. This process introduces a pair of special tokens, <|answer|> and <|use_tool|>, as shown in Figure 1(e) and (g). If the model chooses to provide a direct answer, it outputs <|answer|> first; if it chooses function calling, it outputs <|use_tool|> first. This classification task forces the model to make a decision based on the user query and provided functions before delving into the details of function calling, thereby enhancing the stability of its output.
The Decision Token also facilitates the creation of non-function-call data from function-called data. To generate non-function-call data, consider an example where the original data involves three functions: func_A, func_B, and func_C. Based on the user query, func_A is helpful and thus called in the original data. By assuming that func_B and func_C are not helpful, we can create non-function-call data by removing func_A. The remaining functions, func_B and func_C, will not trigger any function calling for the user query. This allows us to easily obtain non-function-call data. Previously, generating non-function-call data for training was challenging because it required specific LLM responses for non-function-call cases. However, with the Decision Token, we can train the model to output only <|answer|> in non-function-call cases. During inference, this is not an issue because the model will continue to provide an appropriate response after <|answer|>.

### v2

@@ 1 Introduction
Figure 1: An illustration of prompt templates used for function calling and instruction following in LLMs. During training, LLMs are given conditional prompts (shown on the left) and tasked with generating corresponding text completions (shown on the right). When a function call is required, the model generates structured function calls in the form of a list of functions, where each function is specified with its arguments in the format func_name(arg1=value1, ...). Special tokens, including <s>, <|im_start|>, <|im_end|>, <|answer|>, and <|use_tool|>, are each represented by a single token after tokenization. For more details, refer to Section 3.1.
@@ 3.1 Prompt Templates for Function Calling and Instruction Following
We employ a tuning-based approach to enable both function-calling and instruction-following capabilities in our LLMs. This involves fine-tuning pretrained base models using prompt templates based on the Chat Markup Language (ChatML), a widely adopted format introduced by OpenAI.
Two main strategies for incorporating function descriptions into prompts are explored: (1) introducing a dedicated role, such as tools, to represent function descriptions in JSON format (Figure 1(b)); and (2) embedding function descriptions alongside usage instructions within the system role (Figure 1(c)). In the latter strategy, both instruction-following and function-calling are guided by the system prompt.
During training, the LLMs are provided with conditional prompts as described above and are tasked with generating appropriate text completions. Based on the context, the fine-tuned model dynamically decides whether to respond directly or invoke functions. If no relevant functions are available, the model directly answers the query (Figure 1(d)). Otherwise, if function calls are needed, the model generates structured function calls in the form of a list of functions (Figure 1(f)).
In the experiments, we investigated the performance comparison of different conditional prompts and the use of training data across various metrics for instruction-following and function-calling capabilities, as discussed in Section 4.2.
@@ 3.2 Decision Token
To address this, we propose the novel Decision Token mechanism. LLMs generate responses through next-token prediction, where each step involves a classification task to select the next token. The Decision Token concept leverages the fact that each token prediction is essentially a classification. By introducing a pair of special tokens, the model can predict a binary classification that determines whether to answer the query directly or invoke function calls before generating a detailed response or function calls, respectively. Specifically, this process introduces a pair of special tokens, <|answer|> and <|use_tool|>, as shown in Figure 1(e) and (g). If the model chooses to provide a direct answer, it outputs <|answer|> first; if it chooses function calling, it outputs <|use_tool|> first. This classification task forces the model to make a decision based on the user query and provided functions before delving into the details, thereby enhancing the stability of its output.
The Decision Token also facilitates the creation of non-function-call data from function-called data. To generate non-function-call data, consider an example where the original data involves three functions: func_A, func_B, and func_C. Based on the user query, func_A is helpful and thus called in the original data point. By assuming that func_B and func_C are not helpful, we can create non-function-call data by removing func_A as input. With only func_B and func_C as the remaining functions, function calling should not be triggered from user query and a direct answer should be provided. This allows us to easily obtain non-function-call data. Previously, generating non-function-call data for training was challenging because it required specific LLM responses for non-function-call cases. However, with the Decision Token, we can train the model to output only <|answer|> in non-function-call cases. During inference, this is not an issue because the model will continue to provide an appropriate response after <|answer|>.